OpenStudy (anonymous):

I am in desperated need for some algebra 2 help!! I will fan and medal anyone who helps me I take a test tomorrow and I have no idea what i'm doing!!

OpenStudy (anonymous):

What is the solution to the rational equation x over 2x minus 1 plus 1 over 4 equals 2 over 2x minus 1 ? x = 9 over 2 x = 7 over 2 x = 3 over 2 x = 7 over 6

OpenStudy (johnweldon1993):

\[\large \frac{x}{2x - 1} + \frac{1}{4} = \frac{2}{2x - 1}\] like that?

OpenStudy (anonymous):

Yes, that's what it looks like.

OpenStudy (anonymous):

Are you able to help me? @johnweldon1993

OpenStudy (johnweldon1993):

I'd certainly like to think so... Multiply everything by \(\large 2x - 1\) \[\large 2x - 1 \times \frac{x}{2x - 1} + 2x-1 \times \frac{1}{4} = 2x - 1\frac{2}{2x - 1}\] Notice \[\large \cancel{2x - 1} \times \frac{x}{\cancel{2x - 1}} + 2x-1 \times \frac{1}{4} = \cancel{2x - 1}\frac{2}{\cancel{2x - 1}}\] so all we have left is \[\large x + \frac{2x-1}{4} = 2\] look better? what would you do next?

OpenStudy (anonymous):

yes and umm i'm not really sure this stuff confuses me.. @johnweldon1993

OpenStudy (johnweldon1993):

Well, did what I do confuse you?

OpenStudy (anonymous):

No, I got that part I just don't know what to do next.

OpenStudy (johnweldon1993):

Okay good, well that was the worst part... now \[\large x + \frac{2x - 1}{4} = 2\] If we multiply everything by 4...we would have \[\large 4x + \cancel{4}\times \frac{2x -1}{\cancel{4}} = 8\] \[\large 4x + 2x - 1 = 8\] what next?

OpenStudy (anonymous):

Then you'd combine like terms

OpenStudy (johnweldon1993):

Correct....which would leave you with...?

OpenStudy (anonymous):

6x-1=8

OpenStudy (johnweldon1993):

Now just solve for 'x'

OpenStudy (anonymous):

x=9/6

OpenStudy (anonymous):

which then reduces to 3/2

OpenStudy (johnweldon1993):

And that is your answer! :)

OpenStudy (anonymous):

Can you help me with come more? I'd really appreciate it!

OpenStudy (johnweldon1993):

Of course :)

OpenStudy (anonymous):

Want me to open a new question or just do it on this one?

OpenStudy (johnweldon1993):

Hmm just do it here, no need to open a new one :)

OpenStudy (anonymous):

What is the simplified form of the quantity 4 x squared minus 25 over the quantity 2 x plus 5 ? 2x - 5, with the restriction x ≠ -five over 2 2x + 5, with the restriction x ≠ five over 2 2x - 5, with the restriction x ≠ five over 2 2x + 5, with the restriction x ≠ -five over 2

OpenStudy (johnweldon1993):

\[\large \frac{4x^2 - 25}{2x + 5}\] So right here....a restriction would be when the denominator = 0...because when the denominator = 0 we have an undefined function...so when does \[\large 2x + 5 = 0\]?

OpenStudy (anonymous):

you'd get x=-5/2

OpenStudy (johnweldon1993):

Correct! That is your restriction

OpenStudy (anonymous):

Ohhhh!

OpenStudy (johnweldon1993):

Now, we just need to simplify the expression

OpenStudy (johnweldon1993):

\[\large \frac{4x^2 - 25}{2x + 5}\] So...how can we simplify that top part?

OpenStudy (anonymous):

(4x-5)(x+5)?

OpenStudy (johnweldon1993):

Very close! however when you expand that back out you get 4x^2 + 20x - 5x - 25 4x^2 + 15x - 25 so not quite....any other thoughts?

OpenStudy (anonymous):

umm no not really I thought that was the right way to simplify that

OpenStudy (johnweldon1993):

well remember,4 has 2 different ways to factor...we can have 4 and 1.....OR you can do 2 and 2 right?

OpenStudy (anonymous):

Yes

OpenStudy (johnweldon1993):

Soooo, lets try out (2x + 5)(2x - 5) expand that out..what do we get?

OpenStudy (anonymous):

4x^2-10x+10x-25

OpenStudy (johnweldon1993):

mmhmm, and combine like terms... we get back to the original 4x^2 - 25 right?

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

It's always good to check if you can get back to the start... so now we have \[\large \frac{(2x - 5)(2x + 5)}{2x + 5}\] Well notice... \[\large \frac{(2x - 5)\cancel{(2x + 5)}}{\cancel{2x + 5}}\] so all we're left with is \[\large 2x - 5\]

OpenStudy (anonymous):

ahhh its all making sense now.

OpenStudy (johnweldon1993):

Good :) remember, ask anything that doesn't make sense. Might as well ask while I'm here :P

OpenStudy (anonymous):

What is the solution to the rational equation x over x squared minus 9 plus 1 over x minus 3 equals 1 over 4x minus 12 ?

OpenStudy (anonymous):

You have been a ton of help I really appreciate it!

OpenStudy (johnweldon1993):

\[\large \frac{x}{x^2 - 9} + \frac{1}{x - 3} = \frac{1}{4x - 12}\] any idea where to start with this one?

OpenStudy (anonymous):

Well I would start by I guess trying to like cancel out the bottom terms

OpenStudy (johnweldon1993):

Right....go ahead and let me know when you get stuck :)

OpenStudy (anonymous):

\[\frac{ x }{ x+3 }+\frac{ 1 }{ x-3 }=\frac{ 1 }{4}\] thats what I got so far

OpenStudy (johnweldon1993):

wait...where did that come from?

OpenStudy (anonymous):